# Multi-period Stochastic Network Design for Combined Natural Gas and Hydrogen Distribution

### v3

@@ 5 Computational Experiments
This section evaluates the computational efficiency of our progressive-hedging-based matheuristic, focusing on the speed and scalability of the method. We compare our two-stage stochastic programming solution (SP) with the expected value solution (EV), which is computed by taking the expected values of stochastic parameters and evaluating its solution for each scenario. This help us to measure the Value of Stochastic Solution (VSS), defined as VSS = EV - SP, the value of taking uncertainty into account in the decision-making processes. In Section 5.1, we introduce the benchmark instances and computational parameters. In Section 5.2, we give an overview of the efficiency of our developed methods, where we also briefly discuss an alternative solution method based on Benders decomposition. In Section 5.3, we test the impact of the bundle size on our matheuristic. In Section 5.4, we assess the influence of various uncertainty levels on our matheuristic. In Section 5.5, we discuss the difference in solution structure when accounting for uncertainty compared to taking the expected net supply. The case study is presented in Section 6.
@@ 5.1 Benchmark Instances and Computational Parameters
All the experiments are performed using an AMD 7763 CPU (with 16 cores @ 2.45 GHz) with 32GB RAM. All algorithms are executed in C++20 in combination with Gurobi 10.0.1. We generate $20$ instances for some combinations of $N\in\{8,10,12,15,20\}$, $T\in\{8,10,15\}$, and $W\in\{60,120,240\}$. The number of arcs, $A$, varies per each instance by the instance generator. The average number of arcs for these instances is reported later in our computational results table. Each of these instances is solved by our progressive-hedging-based matheuristic (PHM) and Gurobi (GRB). We also obtain the expected value solution (EV) by solving a single-scenario expected value problem and then simulating the performance of this solution over all scenarios in $\Omega$. For PHM, we apply a MIP gap of $10\%$ for the Initialization and Update steps of Algorithm 4. We further apply a time limit of $1$ hour and a MIP gap of $1\%$ for the Solution step. We set $p_{\textsc{h}}$ and $p_{\textsc{e}}$ to $0.2$ and $1$, respectively. We multiply these probabilities by $0.97$ in each iteration of the algorithm for earlier convergence. We select the bundle size of PHM as $6$ for the instances with $W=60$, and $8$ for $W=\{120,240\}$ according to the preliminary experiments. In Section 5.3 we provide further insights on the bundle size. We solve the base model (1)-(17) via GRB. The EV solution is obtained by taking the average of all stochastic parameters to create an average scenario for each instance. Both GRB and EV have a $3$-hour time limit. The GRB MIP gap is set as $1\%$. We believe this is a competitive setting to test for GRB, since without these limits it either takes too much time to solve the problem or the chances of going out of memory are too high.
@@ 5.2 Performance of the Matheuristic
In this section, we compare our progressive-hedging-based matheuristic (denoted as PHM) with the MIP solution of the base model (denoted as GRB), and the expected value solution (denoted by EV). A summary of the average performance over all instances is provided in Table 1.
Table 1: Results of the relative performance of PHM and EV compared to GRB
@@
[…]
$N$ $A$ $T$ $W$$\#_{\textsc{GRB}}$$\#_{\textsc{PHM}}$ min avg min avg min avg min avg
@@
Each row in Table 1 denotes the performance over 20 instances. Column $A$ denotes the average number of arcs over the $20$ instances. The columns #$\#_{\textsc{GRB}}$ and #$\#_{\textsc{PHM}}$ are the number of instances that are solved without any memory issues. We do not report the number of solutions for the EV, since it stays within the memory limit for all instances. The following columns are relative comparisons. First, we compare PHM relative to GRB both in objective value (Gap (%)) and time (Time (%)). These values are derived from the instances that are solved within the memory limits both with GRB and PHM. In case #$\#_{\textsc{GRB}}=0$, we compare the solution time of PHM with the limit of $3$ hours. The same information is provided for the EV solution, however here, we compared the objective to the best-found solution by either GRB or PHM.
@@
We also make comparisons with the EV. In the column “Gap” of EV, the EV solution is compared with the best solution found via GRB and PHM. Note for the cases that GRB cannot solve any instance, this is always the solution provided by PHM. The solution time of EV is the percentage of GRB time, except for the cases #$\#_{\textsc{GRB}}=0$ where we compare the EV time by $3$ hours. While EV works fast on small instances, we observe an overall gap of $10-20\%$ in its solution quality. Moreover, there is an increase in its time compared to GRB on larger instances, since GRB starts hitting the $3$-hour limit more and more. PHM has similar time ratios on the same larger instances while having better results on the gap. In order to better understand the effect of computational limits on performance comparisons, we conducted additional tests on instances with $N=10$, $T=8$, and $W=60$, but without any time limits and MIP gaps, and using 70GB of RAM for both GRB and PHM. For these specific instances, both the gap and the time required decreased for PHM on average compared to GRB; the gap reduced from $2.52\%$ to $1.46\%$, and the time ratio from $5.65\%$ to $1.50\%$. These results support that imposing limits is beneficial for maintaining GRB’s efficiency compared to PHM.
@@
It is important to note that for the instances with $N=20$, PHM improves the EV solution by about $10\%$ while spending half of the EV solution time. This is remarkable since while the EV is a $1$-scenario problem, the PHM deals with thirty $8$-scenario problems on $W=240$. The algorithm’s success can be attributed not only to the high MIP gap but also to the variable fixing through heuristic constraints, which enhances efficiency. Each iteration is solved as fast as possible, and at the end of each iteration, we fix decision variables as much as possible while keeping overall solution quality high. Only in one extreme case on $N=15$, $T=10$, and $W=240$, we observe that the EV is significantly better (due to PHM not converging quickly enough) as it provides a solution with a $-53.25\%$ gap. However, even when this is included, on average these instances are improved by $11.57\%$ in terms of the solution quality by PHM.
Finally, we address the performance of Benders decomposition based methodology. Initially, we explored the use of Benders decomposition, an approach for stochastic network design problems (Crainic et al., 2021; Rahmaniani et al., 2018, 2024), see B for details. Preliminary experiments with Benders decomposition, however, indicated significant limitations in terms of computational efficiency for our problem. In order to handle with these inefficiencies, Crainic et al. (2021) accelerate a partial Benders decomposition with some strategies such as adding cuts, MIP gaps and time limits for a stochastic network design study, resulting in optimality gaps about $1-5\%$ for problems mostly sized at $N=10$ with no multi-period setting. Similar MIP gaps, and time limits are also implemented in the other Benders decomposition based network design literature (Rahmaniani et al., 2018, 2024), resulting in sub-optimal solutions. Because of this, as the problem complexity increases, particularly with multi-period settings, it is usual to apply heuristics rather than exact solution methods (Boland et al., 2017; Jiang et al., 2021; Fragkos et al., 2021). Overall, consistent with these approaches, we believe our algorithm offers a balance of solution quality and computational time suited for a complex stochastic network design setting, while providing optimality gaps on par with those achieved with Benders decomposition. Notably, while Benders decomposition is theoretically exact, its practical implementation involving MIP gaps and time limits, making its degree of exactness comparable to our heuristic approach, based on progressive hedging which is also an exact method on the base version (Rockafellar and Wets, 1991).
@@ 5.3 Impact of Scenario Count per Bundle on Solution
This subsection analyzes how varying the number of scenarios per bundle affects the solution quality of the PHM. We randomize various instances by using the same parameter settings as before. We test various numbers of scenarios per bundle, i.e. $|\xi|\in\{1,2,3,4,6,8,12,24\}$ to randomly assign the scenarios into bundles. We report their solution in terms of the quality and the solution time in Figure 1, where Figure 1(a) and Figure 1(b) provide the individual and the average behavior, respectively.
(a) Individual solution hexagon map of 200 instances with $(N,T,W)=(10,8,120)$ on different bundling; $|\xi|\in\{1,8,12\}$
1111222233334444666688881212121224242424772772772772774774774774776776776776$|\xi|$Obj100100100100200200200200300300300300400400400400Time (sec)
(b) Average solution of 200 instances with $(N,T,W)=(5,8,120)$ or $(N,T,W)=(8,8,120)$ with $|\xi|\in\{1,2,3,4,6,8,12,24\}$
Figure 1: The effects of $|\xi|$ in solution quality and computational time
@@
A similar behavior is seen in Figure 1(b). The figure provides the average objective value and corresponding solution times of a total of $200$ instances with either $(N,T,W)=(5,8,120)$ or $(N,T,W)=(8,8,120)$. Each of these instances is solved for various randomized bundles with $|\xi|\in\{1,2,3,4,6,8,12,\allowbreak 24\}$. We observe that the trade-off observed in Figure 1(a) persists, except for $|\xi|\in\{2,3\}$. This trade-off is due to using our soft variable fixing approach, which incorporates heuristic constraints when a certain percentage of instances share common pipeline arcs. As $|\xi|$ increases, even a few bundles sharing the same arc can trigger the heuristic constraints. In the highest value of $|\xi|=24$, heuristic constraints are introduced even if only one bundle contains a particular arc. Because when $W=120$, $p_{\textsc{h}}=20\%$ of the solutions consist of exactly one bundle with $|\xi|=24$ scenarios. Thus, with higher values of $|\xi|$, the incorporation of heuristic constraints becomes more frequent. While this restriction on the solution space indeed tends to increase the objective value, it also leads to a decrease in the solution time.
For $|\xi|=\{1,4,6,8,12,24\}$, the decision-maker may select the trade-off between solution quality and solution time. Regardless of the chosen trade-off, our method consistently exhibits small objective gaps across all values of $|\xi|$, underscoring its robustness and efficacy. We select either $|\xi|=6$ or $|\xi|=8$ for the computational experiment instances. This choice is influenced by the observation that the average objective gap appears to be relatively insensitive to the number of scenarios in our particular setting, and for $6$ or more values the solution time decreases sharply. However, the objective gap range highly depends on the instance parameters selected. For instance, scenarios characterized by higher uncertainty in supply/demand and high penalty costs may exhibit more substantial differences in objective value. In this case, it might be sensible to decrease $|\xi|$ to find a better balance between the objective function value and the solution time. Similarly, for larger instances requiring extended computational time, it may be advantageous to increase the number of scenarios $|\xi|$ even further, aiming to find some feasible solutions in reasonable computational time. This is particularly important because as computational time increases, the likelihood of encountering memory issues also rises.
@@ 5.4 Impact of the Uncertainty on Solution Quality
In this section, we examine the impact of the uncertainty level on the performance of our progressive-hedging-based matheuristic (PHM) compared to the Gurobi (GRB). We analyze instances with $N=12$ and $A=134$, on low, normal, and high uncertainty in scenarios to explore their influence on gap and solution time. These solutions are executed in C++20 in combination with Gurobi 11.0.0. Normal uncertainty instances are the same as those in Section 5.2. Low uncertainty instances are created similar to the normal ones, but we remove the uncertainty in supply and intermediate locations for hydrogen (see A for details). High uncertainty instances are introduced by adjusting the range of variability parameters as described in A: $[G_{L},G_{U}]=[1.2,1.3]$, $[H_{L},H_{U}]=[1.75,2.25]$, and setting $r$ to be uniformly randomized within $[1.01,1.1]$ for Eq. (46). The results are presented in Table 2.
Table 2: Results of the relative performance of PHM and EV compared to GRB for $N=12$ and $A=134$
@@
[…]
Var $T$ $W$$\#_{\textsc{GRB}}$$\#_{\textsc{PHM}}$ min avg min avg min avg min avg
@@ 5.5 Timing of the Transition in the Hydrogen Economy
In this section, we investigate the dynamics of transitioning from a natural gas network to one primarily based on hydrogen. We conduct 100 instances on each of the parameter sets $(N,T,W)=(8,10,60)$, representing a smaller network with fewer nodes and potential connections, and $(N,T,W)=(12,10,60)$, illustrating a larger, more complex network structure. While our instances are designed to evaluate computational performance, they are derived to reflect real-world conditions and closely align with the parameters of actual case studies in the hydrogen economy, see A for details, making them practical for our analysis in this section. We compare the average proportion of network arcs allocated to hydrogen transportation in each period. All these instances are solved via PHM and EV. Figure 2 illustrates the evolution of hydrogen implementation over time, where the y-axis represents the fraction of arcs being designated for hydrogen transport.
[…]
Figure 2: Hydrogen implementation over time
@@
2222444466668888000.50.50.50.511111.51.51.51.52222Time PeriodFlow Cost RatioPHMEV
(a) Average flow cost per period
22224444666688880011112222333344445555Time PeriodPenalty Cost RatioPHMEV
(b) Average penalty cost per period
Figure 3: Expected flow and penalty cost per period of PHM and EV

### v4

@@ 5 Computational Experiments
This section presents a computational study to evaluate our progressive-hedging-based matheuristic (PHM). We assess its performance against two key benchmarks: the direct solution using Gurobi (GRB) and the solution from the deterministic expected value (EV) problem. The comparison with GRB serves to validate our matheuristic’s computational efficiency and its ability to find high-quality solutions for large-scale instances where exact methods may struggle. The comparison with the EV solution, in turn, demonstrates the value of the stochastic approach and allows us to analyze the structural differences in the resulting network designs when uncertainty is explicitly considered.
In Section 5.1, we introduce the benchmark instances and computational parameters. In Section 5.2, we give an overview of the efficiency of our developed methods. In Section 5.3, we test the impact of the bundle size on our matheuristic. In Section 5.4, we assess the influence of various uncertainty levels on our matheuristic. In Section 5.5, we discuss the difference in solution structure when accounting for uncertainty compared to taking the expected net supply. The case study is presented in Section 6.
@@ 5.1 Benchmark Instances and Computational Parameters
All the experiments are performed using an AMD 7763 CPU (with 16 cores @ 2.45 GHz) with 32GB RAM. All algorithms are executed in C++20 in combination with Gurobi 10.0.1. We generate $20$ instances for some combinations of $N\in\{8,10,12,15,20\}$, $T\in\{8,10,15\}$, and $W\in\{60,120,240\}$. These scenario counts were chosen to balance uncertainty representation and computational tractability. They also align with common practices in stochastic network design, such as Crainic et al. (2021), who use 9625696–256 scenarios for comparable problem sizes. The values of $W$ are also selected for their high divisibility, enabling testing of various bundle sizes. Based on these tests in Section 5.3, we use bundle sizes of $6$ for the case $W=60$, and $8$ for the cases $W=\{120,240\}$. We consider a setting with two commodities: one representing natural gas and one representing hydrogen. These are treated as distinct, homogeneous energy types without differentiation by quality grade. The number of arcs, $A$, varies per each instance by the instance generator. The average number of arcs for these instances is reported later in our computational results table.
Each of these instances is solved by our progressive-hedging-based matheuristic (PHM) and Gurobi (GRB). We also obtain the expected value solution (EV) by solving a single-scenario expected value problem and then simulating the performance of this solution over all scenarios in $\Omega$. For PHM, we apply a MIP gap of $10\%$ for the Initialization and Update steps of Algorithm 4. We further apply a time limit of $1$ hour and a MIP gap of $1\%$ for the Solution step. We set $p_{\textsc{h}}$ and $p_{\textsc{e}}$ to $0.2$ and $1$, respectively. We multiply these probabilities by $0.97$ in each iteration of the algorithm for earlier convergence. We solve the base model (1)-(17) via GRB. The EV solution is obtained by taking the average of all stochastic parameters to create an average scenario for each instance. Both GRB and EV have a $3$-hour time limit. The GRB MIP gap is set as $1\%$. We believe this is a competitive setting to test for GRB, since without these limits it either takes too much time to solve the problem or the chances of going out of memory are too high.
@@ 5.2 Performance of the Matheuristic
In this section, we compare our progressive-hedging-based matheuristic (denoted as PHM) with the MIP solution of the base model (denoted as GRB), and the expected value solution (denoted by EV). Initially, we explored using Benders decomposition; however, preliminary experiments indicated this approach was not computationally competitive for our problem. Its details are given in B. A summary of the average performance over all instances is provided in Table 2.
Table 2: Solution quality (Gap %) and time (Time %) comparison for the matheuristic (PHM) and Expected Value (EV) solutions. All values are relative to a benchmark solution (either GRB or PHM), with precise definitions provided in the subsequent text.
@@
Gap ($\%$) Time ($\%$) Gap ($\%$) Time ($\%$)
NN AA $T$ $W$ $_{\textsc{GRB}}$ $_{\textsc{PHM}}$ min avg min avg min avg min avg
@@
Each row in Table 2 denotes the performance over 20 instances. Column $A$ denotes the average number of arcs over the $20$ instances. The columns $_{\textsc{GRB}}$ and $_{\textsc{PHM}}$ are the number of instances that are solved without any memory issues. We do not report the number of solutions for the EV, since it stays within the memory limit for all instances. The remaining columns compare PHM and EV with GRB in terms of objective value (Gap %) and solution time (Time %). Gaps are computed using the best-found feasible solutions, using only the instances solved by both methods without memory issues. For PHM, the gap is the percentage difference from the GRB solution. For EV, the gap is the percentage difference to the best solution found by either GRB or PHM; thus in cases where GRB fails on all instances, the PHM solution fully serves as the gap benchmark. “Time (%)” columns report the solution time relative to the GRB run time. In cases where $_{\textsc{GRB}}=0$, the time is instead compared against the 3-hour time limit.
@@
We also make comparisons with the EV. In the column “Gap” of EV, the EV solution is compared with the best solution found via GRB and PHM. Note for the cases that GRB cannot solve any instance, this is always the solution provided by PHM. The solution time of EV is the percentage of GRB time, except for the cases $_{\textsc{GRB}}=0$ where we compare the EV time by $3$ hours. While EV works fast on small instances, we observe an overall gap of $10-20\%$ in its solution quality. Moreover, there is an increase in its time compared to GRB on larger instances, since GRB starts hitting the $3$-hour limit more and more. PHM has similar time ratios on the same larger instances while having better results on the gap. In order to better understand the effect of computational limits on performance comparisons, we conducted additional tests on instances with $N=10$, $T=8$, and $W=60$, but without any time limits and MIP gaps, and using 70GB of RAM for both GRB and PHM. For these specific instances, both the gap and the time required decreased for PHM on average compared to GRB; the gap reduced from $2.52\%$ to $1.46\%$, and the time ratio from $5.65\%$ to $1.50\%$. These results support that imposing limits is beneficial for maintaining GRB’s efficiency compared to PHM.
@@
It is important to note that for the instances with $N=20$, PHM improves the EV solution by about $10\%$ while spending half of the EV solution time. This is remarkable since while the EV is a $1$-scenario problem, the PHM deals with thirty $8$-scenario problems on $W=240$. The algorithm’s success can be attributed not only to the high MIP gap but also to the aggregate soft fixing through heuristic constraints, which enhances efficiency. Each iteration is solved as fast as possible, and at the end of each iteration, we fix decision variables as much as possible while keeping overall solution quality high. Only in one extreme case on $N=15$, $T=10$, and $W=240$, we observe that the EV is significantly better (due to PHM not converging quickly enough) as it provides a solution with a $53.25\%$ gap. However, even when this is included, on average these instances are improved by $11.57\%$ in terms of the solution quality by PHM.
@@ 5.3 Impact of Scenario Count per Bundle on Solution
This subsection analyzes how varying the number of scenarios per bundle affects the solution quality of the PHM. We randomize various instances by using the same parameter settings as before. We test various numbers of scenarios per bundle, i.e. $|\xi|\in\{1,2,3,4,6,8,12,24\}$ to randomly assign the scenarios into bundles, once for each combination of instance and $|\xi|$. We report their solution in terms of the quality and the solution time in Figure 1, where Figure 1(a) and Figure 1(b) provide the individual and the average behavior, respectively.
(a) Hexbin map of individual solutions for 200 instances with $(N,T,W)=(10,8,120)$ under different bundling sizes ($|\xi|\in{1,8,12}$). Color intensity corresponds to the concentration of solutions.
[…]
(b) Average objective (left axis) and solution time (right axis) over a combined set of 200 instances (100 of size $(N,T,W)=(5,8,120)$ and 100 of size $(N,T,W)=(8,8,120)$) for varying bundle sizes $|\xi|\in\{1,2,3,4,6,8,12,24\}$
Figure 1: Impact of the number of scenarios ($|\xi|$) on solution quality and computation time across different instance sets.
@@
A similar behavior is seen in Figure 1(b). The figure provides the average objective value and corresponding solution times of a total of $200$ instances with either $(N,T,W)=(5,8,120)$ or $(N,T,W)=(8,8,120)$. Each of these instances is solved for various randomized bundles with $|\xi|\in\{1,2,3,4,6,8,12,\allowbreak 24\}$. We observe that the trade-off observed in Figure 1(a) persists, except for $|\xi|\in\{2,3\}$. This trade-off is due to using our soft fixing approach, which incorporates heuristic constraints when a certain percentage of instances share common pipeline arcs. As $|\xi|$ increases, even a few bundles sharing the same arc can trigger the heuristic constraints. In the highest value of $|\xi|=24$, heuristic constraints are introduced even if only one bundle contains a particular arc. Because when $W=120$, $p_{\textsc{h}}=20\%$ of the solutions consist of exactly one bundle with $|\xi|=24$ scenarios. Thus, with higher values of $|\xi|$, the incorporation of heuristic constraints becomes more frequent. While this restriction on the solution space indeed tends to increase the objective value, it also leads to a decrease in the solution time.
For $|\xi|=\{1,4,6,8,12,24\}$, the decision-maker may select the trade-off between solution quality and solution time. Regardless of the chosen trade-off, our method consistently exhibits small objective gaps across all values of $|\xi|$, underscoring its robustness and efficacy. To justify our focus on average values, we examined the distribution of objective values across the 200 randomized runs. The results show limited variability: the interquartile range remains below $0.75\%$ of the average objective for all bundle sizes —except $|\xi|=24$ which showed a slightly increased range of $1.36\%$— indicating that random scenario grouping introduces no significant variance.
As introduced at the beginning of this section, we select either $|\xi|=6$ or $|\xi|=8$ for our computational experiments. This choice is influenced by the observation that the average objective gap appears to be relatively insensitive to the number of scenarios in our particular setting, and for $6$ or more values the solution time decreases sharply. However, the objective gap range highly depends on the instance parameters selected. For instance, scenarios characterized by higher uncertainty in supply/demand and high penalty costs may exhibit more substantial differences in objective value. In this case, it might be sensible to decrease $|\xi|$ to find a better balance between the objective function value and the solution time. Similarly, for larger instances requiring extended computational time, it may be advantageous to increase the number of scenarios $|\xi|$ even further, aiming to find some feasible solutions in reasonable computational time. This is particularly important because as computational time increases, the likelihood of encountering memory issues also rises.
@@ 5.4 Impact of the Uncertainty on Solution Quality
In this section, we examine the impact of the uncertainty level on the performance of our progressive-hedging-based matheuristic (PHM) compared to the Gurobi (GRB). We analyze instances with $N=12$ and $A=134$, on low, normal, and high uncertainty in scenarios to explore their influence on gap and solution time. These solutions are executed in C++20 in combination with Gurobi 11.0.0. Normal uncertainty instances are the same as those in Section 5.2. Low uncertainty instances are created similar to the normal ones, but we remove the uncertainty in supply and intermediate locations for hydrogen (see A for details). High uncertainty instances are introduced by adjusting the range of variability parameters as described in A: $G_{L},G_{U}]=[1.2,1.3]$, $H_{L},H_{U}]=[1.75,2.25]$, and setting $r$ to be uniformly randomized within $[1.01,1.1]$ for Eq. (49). The results are presented in Table 3.
Table 3: Performance of PHM and EV relative to GRB for instances with $N=12$ and $A=134$, analyzed across Low, Normal, and High uncertainty (Var) settings.
@@
Gap ($\%$) Time ($\%$) Gap ($\%$) Time ($\%$)
Var $T$ $W$ $_{\textsc{GRB}}$ $_{\textsc{PHM}}$ min avg min avg min avg min avg
@@ 5.5 Timing of the Transition in the Hydrogen Economy
In this section, we investigate the dynamics of transitioning from a natural gas network to one primarily based on hydrogen. We conduct experiments using 100 instances on each of the parameter sets $(N,T,W)=(8,10,60)$, representing a smaller network with fewer nodes and potential connections, and $(N,T,W)=(12,10,60)$, illustrating a larger, more complex network structure. While our instances are designed to evaluate computational performance, they are derived to reflect real-world conditions and closely align with the parameters of actual case studies in the hydrogen economy, see A for details, making them practical for our analysis in this section. We compare the average proportion of network arcs allocated to hydrogen transportation in each period. All these instances are solved via PHM and EV. Figure 2 illustrates the evolution of hydrogen implementation over time, where the y-axis represents the fraction of arcs being designated for hydrogen transport.
22446688101000.20.20.40.40.60.6Time PeriodHydrogen RatioN=8N=8, PHMN=8N=8, EVN=12N=12, PHMN=12N=12, EV
Figure 2: Evolution of the hydrogen network over time, comparing the proportion of network arcs designated for hydrogen for PHM and EV solutions across two network sizes (N=8 and N=12)
@@
2244668800.50.5111.51.522Time PeriodFlow Cost RatioPHMEV
(a) Flow cost ratio over time.
2244668801122334455Time PeriodPenalty Cost RatioPHMEV
(b) Penalty cost ratio over time.
Figure 3: Average per-period costs for PHM and EV solutions, shown as a ratio relative to the average cost at t=0.